The problem I'm stuck on is as follows: "Find three consecutive whole numbers whose sum is 133." I really thought the formula "x + (x+1) + (x+2) = 133" was the correct one, but, I keep ending up with "x = 43.333333333333333333333.....", which doesn't seem to satisfy the problem. What am I doing wrong?
Answer by richwmiller(17219)

Your solution is correct.
There is something wrong with the problem.
42+43+44=129
43+44+45=132
44+45+46=135

There are not three consecutive whole numbers with the sum of 133.